OpenStudy (anonymous):

Verify each identity. Show work. 1. (sin(x) - cos(x)/cos(x)) +1 = tan(x) 2. (1/tan(x)) - (sin(x)/cos(x) -1) = -csc(x) 3. (1+tan^2(x)/sin^2(x)+cos^2(x))=sec^(x) 4. (1/tanB)+tanB=secB cscB

OpenStudy (elisaneedshelp):

I suck at Math Sorry :(

OpenStudy (espex):

You are basically just showing that left hand side is equal to the right. Your first one for example \[\frac{\sin(x) - \cos(x)}{\cos(x)} + 1 = \tan(x)\] can be rewritten as \[\frac{\sin(x)}{\cos(x)} - \frac{\cos(x)}{\cos(x)} + 1 = \tan(x)\] where all you are doing is breaking the fraction. Any time you have something over itself, it equals 1 \[\frac{\sin(x)}{\cos(x)} - 1 + 1 = \tan(x)\] ultimately giving you \[\frac{\sin(x)}{\cos(x)} = \tan(x)\] Which is true. Just carry on in the same manner with the rest of them. Good luck.

OpenStudy (anonymous):

@Kainui @danisxl21 @agent0smith @ybarrap

OpenStudy (kainui):

Just start by turning everything into sine and cosine. tanx=sinx/cosx secx=1/cosx cscx=1/sinx sin^2x+cos^2x=1 Those are some good trig identities to get you started. Good luck, if you get stuck at a specific spot, show us your work up to where you get stuck and we can help you on that part.